Ram is four years older than shyam.The sum of their ages is 52 years.Find the present age of shyam.
Answer by chndnbl(31)

Let shyam age be x.
Ram age=x + 4
sum of their ages =52
x + (x + 4) = 52
2x + 4 = 52
2x = 52 - 4
2x = 48
x =48/2
x = 24
Ram age =x + 4 =24 + 4 =28